# Writing – review & editing, Writing – original draft, Visualization, Data curation ,Investigation, Formal analysis, Validation, Software, Methodology, Conceptualization

### v1

@@ Abstract
Machine learning models trained on tabular data are vulnerable to adversarial attacks, even in realistic scenarios iwhereattackers have accessonly tto he model ’soutputs. TResearchers evaluatesuch attacks, by consideringmetrics slikesuccess rate, perturbation magnitude, and query count. However, unlike other data domains, the tabular domain contains complex interdependencies among features, presenting a unique aspect that should be evaluated: the need for the attack to generate coherenctsamples and ensure feature consistency for indistinguishability. Currently, there is no established methodology for evaluating adversarial samples based on these criteria. In this paper, we address this gap by proposing new evaluation criteria tailored for tabular attacks’ quality; we defined anomaly-based framework to assess the distinguishability of adversarial samples and utilize the SHAP explainability technique to identify inconsistencies in the model’s decision-making process caused by adversarial samples. These criteria could form the basis for potential detection methods and be integrated into established evaluation metrics for assessing attack’s quality Additionally, we introduce a novel technique for perturbing dependent features while maintaining coherence and feature consistency within the sample. We compare different attacks’ strategies, examining black-box query-based attacks and transferability-based gradient attacks across four target models. Our experiments, conducted on benchmark tabular datasets, reveal significant differences between the examined attacks’ strategies in terms of the attacker’s risk and effort and the attacks’ quality. The findings provide valuable insights on the strengths, limitations, and trade-offs of various adversarial attacks in the tabular domain, laying a foundation for future research on attacks and defense development.
@@ 1 Introduction
Machine learning (ML) models trained on tabular data are widely deployed in diverse industry sectors including finance (e.g., for computing credit risk) [11, 36], healthcare (e.g., for analyzing administrative claims and patient registry data) [12], and energy (e.g., for predicting energy and electricity consumption) [28, 42]. However, recent research has highlighted their vulnerability to adversarial attacks, in which malicious data samples are used to mislead the model into producing incorrect outputs [2, 7, 35, 19]. For instance, an applicant may slightly increase his reported income (influencing the ‘income level’ feature), crossing a decision threshold and thereby changing the model’s decision from “reject" to “approve." Such adversarial attacks can even be conducted in realistic black-box settings, where the attacker has no knowledge of the model’s internals and can only query the model to receive outputs.
@@
While extensive research has been performed in the image [49, 50, 24, 41], audio [29], text [49, 34, 18], and graph data  [49, 45] domains, the tabular data domain remains relatively underexplored. This paper addresses several key challenges in the realm of adversarial attacks, focusing on tabular data.
@@
Crafting Coherent and Consistent Adversarial Samples: Crafting adversarial samples in the tabular domain presents unique challenges compared to other data types, such as images, where features (pixels) are typically independent and within a fixed range. Tabular data involves diverse feature types and Tabular models often rely on interdependent features, where the values of some features (dependent features) are influenced by others (influencing features). To maintain coherence and avoid detection, perturbations of dependent features must align with the values of other features. For example, ‘BMI’ is derived from ‘height’ and ‘weight’. When modifying these influencing features (i.e., ‘height,’ ‘weight,’ or both) during an attack, it is important to ensure that the ‘BMI’ value remains consistent with the changes made to these features.
While the dependencies of ‘BMI’ are well-defined, there are cases where dependencies are indirect or influenced by unknown features. For example, ‘diabetes risk level’ depends on ‘age,’ ‘BMI,’ ‘family history of diabetes,’ and ‘blood sugar levels.’ In a practical scenario, the influential features may be hidden from the attacker, who might also lack precise knowledge of how the ‘diabetes risk level’ is determined based on them. Therefore, maintaining coherence and consistency within the adversarial sample during the perturbation process becomes particularly challenging.
@@
Evaluation Criteria for Distinguishable Adversarial Samples: Existing evaluation methods for adversarial samples have primarily focused on measuring the perturbation magnitude using $L_{p}$ norms, which are intended to indicate the level of detectability (i.e., the risk that the attack will be exposed). However, in the context of tabular data, simply limiting perturbation magnitudes is insufficient. Effective adversarial attacks must ensure that samples remain coherent, consistent, and in-distribution to evade detection. While previous work has acknowledged these challenges (see  Section 2.1), there is a lack of concrete evaluation methods for quantifying the distinguishability of adversarial samples generated by different attacks in the tabular domain.
@@
Empirical Analysis of Different Attacks’ Strategies: While adversarial attacks have been extensively studied in other domains, the analysis of black-box attacks’ strategies in the tabular domain has been limited. Common black-box attacks’ strategies are query- and transferability-based. In query-based attacks [7], the attacker repeatedly queries the target model, learns its decision boundary from the responses, and optimizes an adversarial sample to mislead it, while in transferability-based attacks [35, 19], the attacker optimizes the adversarial sample to mislead a surrogate model, and then uses this sample to mislead the target model with a single query.
@@
Adversarial attacks are used by both defenders and attackers both of whom may choose different attacks’ strategies based on their goals, resources, and constraints. Defenders who are interested in assessing their models’ vulnerability often use adversarial attacks for automated penetration testing or adversarial training, which involves creating adversarial samples to improve the model’s robustness [10, 21]. Since defenders have full access to their models and can make unlimited queries, query-based attacks are a practical option. In contrast, attackers aiming to minimize queries and avoid detection might prefer transferability-based attacks that use gradient methods on neural networks, as these capture underlying feature interactions necessary for subtle perturbations. Understanding how and whether query-based attacks reflect other attack types, such as transferability-based gradient attacks, is crucial when using them to improve model defenses.
@@
In this paper, we aim to advance adversarial research in the tabular domain by addressing existing gaps in the literature. We compare attacks’ strategies to provide insights for both attackers and defenders and propose a new technique for consistent perturbation of dependent features and concrete evaluation criteria for assessing attacks’ quality. We perform a detailed examination of both black-box query- and transferability-based gradient attacks, tailored to respect tabular distribution constraints.
To address the challenge of crafting coherent and consistent adversarial samples, we introduce a novel technique for perturbing dependent features. This technique uses regression models to correct perturbed features during the attack process, ensuring that they remain consistent with the overall sample (see Section 3).
Additionally, we propose two concrete criteria for the assessment of attacks’ distinguishability and quality (see Section 3), which can also serve as potential detection mechanisms:
  1. 1.
    Anomaly Assessment: Measures the anomaly detection rate of adversarial samples compared to benign samples using isolation forest algorithm and autoencoder.
  2. 2.
    Explainability-Based Analysis: Identifies inconsistencies in the model’s decision-making process caused by adversarial samples, using an anomalous criteria based on SHapley Additive exPlanations (SHAP) [32], which is a commonly used explainability (XAI) technique [3].
Our comparison of attacks’ strategies integrates both traditional metrics of the attacker’s risk (success rate, query count, and $L_{p}$ norms) and effort and the new evaluation criteria introduced in this paper, providing a thorough assessment of adversarial sample quality and attack effectiveness.
Four different target models were used to explore and evaluate the attacks’ strategies (process illustrated in  Fig. 1) on three datasets: the Hateful Users on Twitter [43] and Intensive Care Unit [20] public datasets, which are both benchmark datasets in the tabular domain, and the Video Transmission Quality (VideoTQ) dataset, which is a private dataset. Seven attacks were examined in our evaluation: two well-known query-based attacks (the boundary and HopSkipJump attacks [5, 8] that can be adapted for tabular data [7]) and five transferability-based gradient attacks that varied in terms of their feature selection techniques (a basic attack [35] in which features were randomly selected and four attacks in which features were selected based on their importance [19]).
This work sheds light on the key trade-offs between these attacks for the tabular domain, which can guide the selection of attacks’ strategies and motivate the development of new defense approaches tailored to the unique properties of this domain.
@@
  • We introduce a novel technique for perturbing dependent features while maintaining coherence and ensuring feature consistency with the rest of the features.
  • We propose two evaluation criteria to assess the distinguishability and quality of adversarial samples in the tabular domain, which can also assist the research community in developing more robust and effective attacks and defenses.
  • To the best of our knowledge, this is the first study that thoroughly evaluates the characteristics of and differences between different attacks’ strategies in tabular data domains, which potentially impact defense strategy decisions.
  • The resources of this study (preprocessed data, code, models, and adversarial samples) will be publicly released to promote future research.
The remainder of this paper is organized as follows: In Section 2, we discuss related work. Section 3 introduces our methodological framework, followed by Section 4 and Section 5, which present the experimental setup and provide a thorough analysis of the examined query- and transferability-based attacks’ behavior. Finally, Section 6 and Section 7 conclude with insights for attackers and defenders within tabular data domains.
@@ 2.2 Existing Methods for Evaluating Adversarial Attacks
    Attacker’s risk. The attacker’s risk is assessed by the attack success rate [17, 2], the number of queries made to the target model [22, 47], and the extent to which the adversarial has been distorted, consequently becoming distinguishable. The distortion is typically measured using $L_{p}$ norms such as $L_{0}$, $L_{2}$, and $L_{\infty}$, which quantify the magnitude of perturbations applied to the original samples [17, 27, 14, 2, 35, 19], and can also be evaluated by computing the average error between the original sample and the corresponding adversarial samples[16].
@@ 2.3 Indistinguishability in the Tabular Domain
While most of the metrics used to assess the attacker’s effort and risk are the same for tabular data domains and other domains, the risk of distinguishing an adversarial sample should be measured differently. In the computer vision domain for example, the indication for an indistinguishable adversarial sample is that the human eye is unable to recognize the difference between the adversarial sample and the original sample [46, 17, 39, 27]. It is commonly assessed by measuring the distance between the original and perturbed images using $L_{2}$ norm.
@@
However, in the tabular domain, “human eye" inspection is less meaningful, as only domain experts can accurately examine such differences [2, 7]. Furthermore, when looking at the tabular data samples, a domain expert would “select" a limited number of features to examine and determine whether the sample is real or manipulated by an adversary [2] based on those features. Given the above, the authors proposed to evaluate the distinguishability of a tabular adversarial sample based on the number of “important" features (i.e., features that a domain expert would likely examine) altered in the adversarial attack. Accordingly, an adversarial sample is considered more perceptible as the number of changes in “important" features increases. However, this evaluation method is neither objective nor practical, as it requires a domain expert to manually assess each sample individually. Moreover, it does not consider the fact that altering even a small number of features can cause inconsistency with the rest of the sample, making it distinguishable from benign data distribution and detectable by anomaly-based methods.
@@
Prior research highlighted the significance of both consistency [35, 23] and validity [35, 16] when analyzing the quality of adversarial samples. Validity can be assessed by examining whether each feature value falls within a valid range [2, 35] and by measuring the sample’s deviation from the distribution of benign data [35]. However, no concrete metric has been proposed to quantify the coherence and consistency of adversarial samples. The coherence and consistency of crafted samples are challenging to measure, as it is influenced by correlations between different feature values [35]. To address this gap, we propose two concrete criteria to assess the distinguishability of adversarial samples from benign samples, enabling an indirect evaluation of their coherence and consistency. We assume that adversarial samples lacking coherence or consistency with the benign data distribution are likely to exhibit anomalies, making them both distinguishable and detectable.
@@
3 Methodological Framework for Addressing Challenges in Adversarial Attack and Evaluation within the Tabular Domain
@@
In this section, we introduce the black-box query and transferability attacks evaluated, along with the modifications made to deal with the tabular constraints. We also present our novel technique for perturbing dependent features, as well as new criteria for evaluating the quality of adversarial attacks’ samples. These criteria focus on the distinguishability of the adversarial samples and their impact on the target model’s internal behavior within the tabular domain.
@@
In our evaluation (see Section 4.5), we examine two decision-based attacks: the boundary attack [5] and the HopSkipJump attack [8], which were adjusted to address tabular constraints and evaluated by  [7]. We implemented these attacks by additionally applying constraints based on feature distributions (e.g., limiting values according to statistical information) and making adjustments such as ensuring validity (e.g., binary or categorical constraints) and editability (guided by domain knowledge)
In addition, we performed untargeted transferability-based gradient attacks based on the architecture proposed in [35, 19], which is an advanced attack that aims to craft valid and consistent adversarial samples.
@@
We address the challenge of maintaining coherence and ensuring feature consistency within adversarial samples during perturbation by introducing a novel technique: for each dependent feature, we train a regression model on the benign samples. During the adversarial sample crafting process, these regression models are querying to ensure that perturbations of dependent features remain consistent with the rest of the sample. This method effectively manages both direct and indirect feature dependencies, ensuring the overall sample’s coherence and therefore less distinguishable.
In our implementation of both query- and transferability-based attacks, the attack process includes an additional step in which the dependent features, as determined by domain knowledge, are adjusted based on the regression models.

### v2

@@ Abstract
Machine learning models trained on tabular data are vulnerable to adversarial attacks, even in realistic scenarios where attackers only have access to the model’s outputs. Since tabular data contains complex interdependencies among features, it presents a unique challenge for adversarial samples which must maintain coherence and respect these interdependencies to remain indistinguishable from benign data. Moreover, existing attack evaluation metrics—such as the success rate, perturbation magnitude, and query count—fail to account for this challenge. To address those gaps, we propose a technique for perturbing dependent features while preserving sample coherence. In addition, we introduce Class-Specific Anomaly Detection (CSAD), an effective novel anomaly detection approach, along with concrete metrics for assessing the quality of tabular adversarial attacks. CSAD evaluates adversarial samples relative to their predicted class distribution, rather than a broad benign distribution. It ensures that subtle adversarial perturbations, which may appear coherent in other classes, are correctly identified as anomalies. We integrate SHAP explainability techniques to detect inconsistencies in model decision-making, extending CSAD for SHAP-based anomaly detection. Our evaluation incorporates both anomaly detection rates with SHAP-based assessments to provide a more comprehensive measure of adversarial sample quality. We evaluate various attack strategies, examining black-box query-based and transferability-based gradient attacks across four target models. Experiments on benchmark tabular datasets reveal key differences in the attacker’s risk and effort and attack quality, offering insights into the strengths, limitations, and trade-offs faced by attackers and defenders. Our findings lay the groundwork for future research on adversarial attacks and defense development in the tabular domain.
@@ 1 Introduction
Machine learning (ML) models trained on tabular data are widely deployed in diverse industry sectors including finance (e.g., for computing credit risk) [12, 37], healthcare (e.g., for analyzing administrative claims and patient registry data) [13], and energy (e.g., for predicting energy and electricity consumption) [30, 43]. However, recent research has highlighted their vulnerability to adversarial attacks, in which malicious data samples are used to mislead the model into producing incorrect outputs [2, 8, 36, 21]. For instance, an applicant may slightly increase their reported income (influencing the “income level" feature), crossing a decision threshold and thereby changing the model’s decision from ‘reject’ to ‘approve.’ Such adversarial attacks can even be conducted in realistic black-box settings, where the attacker has no knowledge of the model’s internals and can only query the model to receive outputs.
@@
While extensive research has been performed in the image [50, 52, 26, 42], audio [31], text [50, 35, 20], and graph data  [50, 46] domains, the tabular domain remains relatively underexplored. This paper addresses several key challenges in the realm of adversarial attacks, focusing on tabular data.
Crafting Coherent and Consistent Adversarial Samples: Crafting adversarial samples in the tabular domain presents unique challenges compared to other data types. Unlike domains such as images, audio, and text, which contain homogeneous feature types with well-defined dependencies (spatial in images, temporal in audio, and semantic in text), the tabular domain is characterized by heterogeneous feature types, diverse data constraints, and complex interdependencies. Correlations in tabular data can also be complicated. For instance, in a loan dataset, a high credit score strongly correlates with approval, but for a low credit score, approval depends on other factors like income or debt, making the correlation conditional.
@@
Models trained on tabular data often rely on interdependent features, where the values of some features (dependent features) are influenced by others (influencing features). Therefore, in adversarial attacks, perturbations to dependent features must align with the influencing features to preserve these constraints and interdependencies, maintaining coherence and consistency to avoid detection. For instance, the BMI is derived from height and weight, so when altering these influencing features (such as height or weight) during an attack, it is essential to ensure that the BMI value remains consistent with the changes.
Multiple dependent features further complicate the process of maintaining this alignment. Furthermore, while some dependencies are well-defined (as is the case with the BMI), others may be non-linear, indirect, or influenced by features unknown to the attacker. For instance, the feature “employee performance rating," used in a model predicting company efficiency, depends on factors such as project completion rate, peer feedback, client satisfaction scores, and overtime hours. In practice, the attacker may not have access to all influential features or know exactly how these features contribute to the employee performance rating. These complex, often domain-specific interdependencies and feature correlations make it much harder to maintain consistency and coherence when crafting adversarial samples in the tabular domain.
@@
Evaluation Criteria for Distinguishable Adversarial Samples: Existing evaluation methods for adversarial samples primarily focus on measuring the perturbation magnitude using $L_{p}$ norms, which are intended to indicate the level of detectability (i.e., the risk that the attack will be exposed). However, in the context of tabular data, simply limiting perturbation magnitudes is insufficient. Effective adversarial attacks must ensure that samples remain coherent, consistent, and in-distribution to evade detection. While previous work has acknowledged these challenges (see  Section 2.3), there is a lack of concrete evaluation methods for quantifying the distinguishability of adversarial samples in the tabular domain.
@@
Empirical Analysis of Different Attack Strategies: While adversarial attacks have been extensively studied in other domains, the analysis of black-box attack strategies in the tabular domain has been limited. Query-based and transferability-based are common black-box attack strategies.
In query-based attacks [8], it is assumed that the attacker has unlimited access to the target model’s predictions. By repeatedly querying the model, he attacker is able to iteratively optimize an adversarial example to induce misclassification. Transferability-based attacks [36, 21] operate under a more constrained threat model, where the attacker lacks direct access to the target model’s outputs but is assumed to have access to a large amount of reliable data to train an accurate surrogate model. In this case, the attacker optimizes the adversarial sample to mislead the surrogate model and then uses this sample to mislead the target model with a single query.
@@
Adversarial attacks are used by both defenders and attackers both of whom may choose different attack strategies based on their goals, resources, and constraints. Defenders who are interested in assessing their model’s vulnerability often use adversarial attacks for automated penetration testing or adversarial training, which involves creating adversarial samples to improve the model’s robustness [11, 23]. Since defenders have full access to their model and can make unlimited queries, query-based attacks are a practical option. In contrast, attackers aiming to minimize the number of queries performed and avoid detection, might prefer transferability-based attacks that use gradient methods on neural networks, as they capture underlying feature interactions necessary for subtle perturbations. Understanding how and whether query-based attacks reflect other types of attacks, such as transferability-based gradient attacks, is crucial when using them to improve model defenses.
@@
In this paper, we aim to advance adversarial research in the tabular domain by addressing existing gaps in the literature. We compare the effectiveness of various attack strategies against classification models to provide insights for both attackers and defenders. Additionally, we developed a novel technique for consistent perturbation of dependent features and concrete evaluation criteria for assessing attack quality. We perform a detailed examination of both black-box query- and transferability-based gradient attacks, tailored to respect tabular distribution constraints.
Our proposed technique for perturbing dependent features aims to address the challenge of crafting coherent and consistent adversarial samples. This technique uses regression models to correct dependent features during the attack process, ensuring that they maintain realistic relationships with interdependent features and remain consistent with the values of other features (see Section 3).
To assess the distinguishability and quality of adversarial samples (see Section 3), we introduce Class-Specific Anomaly Detection (CSAD), an approach that measures the anomaly detection rate of adversarial samples relative to the distribution of benign samples from the specific class predicted by the target model. CSAD leverage existing anomaly detection methods to detect anomalies when inconsistencies are present within class-specific distributions. Additionally, we propose complementary criteria that use explainability-based analysis to examine how adversarial samples influence the model’s decision-making process. This analysis uses SHapley Additive exPlanations (SHAP)[33], a widely adopted explainability (XAI) technique[3], to quantify the inconsistencies induced by adversarial perturbations in the model’s internal reasoning.
Our comparison of attack strategies’ performance factors integrates both traditional metrics of the attacker’s risk (success rate, query count, and $L_{p}$ norms) and effort and the new evaluation criteria introduced in this paper (sample anomaly detection and feature importance consistency), providing a thorough assessment of adversarial sample quality and attack effectiveness.
Four different tree-based target models are used to explore and evaluate the attack strategies (process illustrated in  Fig. 1) on three datasets: the Hateful Users on Twitter [44] and Intensive Care Unit [22] public datasets, which are benchmark datasets in the tabular domain, and the Video Transmission Quality (VideoTQ) dataset, which is a proprietary dataset. Seven attacks are examined in our evaluation: two well-known query-based attacks (the boundary and HopSkipJump attacks [6, 9] that can be adapted for tabular data [8]) and five transferability-based gradient attacks that varied in terms of their feature selection techniques (a basic attack [36] in which features were randomly selected and four attacks in which features were selected based on their importance [21]).
This work sheds light on the key trade-offs between these attacks for the tabular domain, which can guide the selection of attack strategies and motivate the development of new defense approaches tailored to the unique properties of this domain.
@@
  • We propose a novel technique for perturbing interdependent features in tabular data, leveraging regression models to preserve coherence and ensure feature consistency, overcoming a key challenge in tabular adversarial attacks.
  • We introduce CSAD, an approach that assesses anomalies relative to class-specific distributions rather than global thresholds, which improve sensitivity to inconsistencies among features.
  • We developed comprehensive dual evaluation criteria based on anomaly detection assessment of the sample values and feature importance, providing a core principled framework to quantify both the quality of tabular adversarial examples.and their impact on model decision-making.
  • To the best of our knowledge, this study is the first to systematically compare adversarial attack strategies’ performance factors in the tabular domain, providing critical insights into their effectiveness, limitations, and corresponding defense strategies.
  • The resources of this study (preprocessed data, code, models, and adversarial samples) will be made publicly available to promote future research and reproducibility.
The remainder of this paper is organized as follows: In Section 2, we discuss related work. Section 3 introduces our methodological framework, followed by Section 4 and Section 5, which present the experimental setup and provide a thorough analysis of the examined query- and transferability-based attack behavior. Finally, Section 6 and Section 7 conclude with insights for attackers and defenders within tabular data domains.
@@ 2.2 Existing Methods for Evaluating Adversarial Attacks
    Attacker’s risk. The attacker’s risk is assessed by the attack success rate [19, 2], the number of queries made to the target model [24, 48], and the extent to which the adversarial sample has been distorted, consequently becoming distinguishable. The distortion is typically measured using $L_{p}$ norms such as $L_{0}$, $L_{2}$, and $L_{\infty}$, which quantify the magnitude of perturbations applied to the original samples [19, 29, 15, 2, 36, 21], and can also be evaluated by computing the average error between the original sample and the corresponding adversarial samples[18].
@@ 2.3 Indistinguishability in the Tabular Domain
While most metrics used to assess the attacker’s effort and risk are consistent across various domains including tabular data, the detectability dimension of adversarial samples should be measured differently for tabular data domains. In the computer vision domain, for example, an adversarial sample is considered to be indistinguishable if the human eye is unable to recognize the difference between the adversarial sample and the original sample [47, 19, 40, 29]. It is commonly assessed by measuring the distance between the original and perturbed images using $L_{2}$ norm.
@@
However, in the tabular domain, “human eye" inspection is less meaningful, as only domain experts can accurately examine such differences [2, 8]. Furthermore, when looking at tabular data samples, a domain expert would “select" a limited number of features to examine and determine whether the sample is real or manipulated by an adversary [2] based on those features. Given the above, the authors of [2, 8] proposed evaluating the distinguishability of a tabular adversarial sample based on the number of “important" features (i.e., features that a domain expert would likely examine) altered in the adversarial attack. Accordingly, an adversarial sample is considered more perceptible as the number of changes in important features increases. However, this evaluation method is neither objective nor practical, as it requires a domain expert to manually assess each sample individually. Moreover, it does not consider the fact that altering even a small number of features can cause inconsistency with the rest of the sample, making it distinguishable from the benign data distribution and detectable by anomaly-based methods.
@@
Prior research highlighted the significance of both consistency [36, 25] and validity [36, 18] when analyzing the quality of adversarial samples. Validity can be assessed by examining whether each feature value falls within a valid range [2, 36] and by measuring the sample’s deviation from the distribution of benign data [36]. However, no concrete metric has been used to quantify the coherence and consistency of adversarial samples, which are influenced by complex correlations between feature values [36].
We address this gap by proposing two criteria that assess the distinguishability of adversarial samples from benign ones using an “anomaly detection rate," which provides an indirect evaluation of their coherence and consistency. We assume that adversarial samples lacking structural coherence or consistency with the benign data distribution are more likely to exhibit anomalies, making them both distinguishable and detectable.
Although there are many well-established anomaly detection techniques for the image domain, not all of them are directly applicable to tabular data. Since some methods fail to fully capture multivariate dependencies, selecting the right approach is crucial. Methods such as convolutional autoencoders (AEs) and generative adversarial networks (GANs) leverage spatial correlations but often struggle to capture the complex dependencies found in tabular data and therefore should be modified for the tabular domain [4]. Similarly, traditional statistical techniques like t-tests and the ANOVA assume normality and independence [16], limiting their ability to detect anomalies in the tabular domain where variables might not follow normal distributions and the relationships between variables are often complex and non-linear. Other simple approaches, such as Z-score thresholding, focus on individual feature deviations but fail to account for multivariate dependencies. Additionally, some commonly used anomaly detection techniques have inherent limitations for tabular data. For instance, PCA-based methods assume linear relationships, limiting their effectiveness in detecting non-linear dependencies in high-dimensional datasets [45].
More suitable approaches for tabular data include AEs, which learn latent representations to detect deviations from expected patterns [45], and the isolation forest (IF) algorithm, which identifies anomalies by isolating points in a partitioned feature space. However, while effective for detecting extreme outliers, this algorithm is less sensitive to subtle structural inconsistencies [51].
@@
3 Methodological Framework for Addressing Challenges in Adversarial Attacks and Evaluation in the Tabular Domain
@@
In this section, we introduce the black-box query- and transferability-based attacks evaluated, along with the modifications made in order to deal with the tabular constraints. We also present our novel technique for perturbing dependent features, as well as two criteria for evaluating the quality of adversarial attack samples. These criteria focus on the distinguishability of the adversarial samples and their impact on the target model’s internal behavior in the tabular domain.
@@
In our evaluation (see Section 4.5), we examined two decision-based attacks: the boundary attack [6] and the HopSkipJump attack [9], which were adjusted to address tabular constraints and evaluated by  [8]. We implemented these attacks by applying additional constraints based on feature distributions (e.g., limiting values according to statistical information) and making adjustments to ensure validity (e.g., [0,1] values for binary features and ordinal values for categorical features) and editability (guided by domain knowledge).
In addition, we performed transferability-based gradient attacks based on the architecture proposed in [36, 21], which is a sophisticated attack designed to craft valid and consistent adversarial samples.
@@
To ensure coherence and maintain feature consistency when crafting adversarial samples, we introduce a novel technique for perturbing dependent features. Specifically, for each dependent feature, we train a regression model on the benign samples, excluding the target feature. This model then predicts the value of the dependent feature based on the values of the other features in the sample.
During the adversarial attack process, when perturbations are applied to features, the regression model corresponding to each dependent feature is queried to predict its value, ensuring that the dependent feature remains consistent with the rest of the sample. This approach ensures that the perturbations respect indirect dependencies between features, thereby maintaining the overall coherence of the sample.
In our implementation of both query- and transferability-based attack strategies, the attack process includes an additional step in which the dependent features, as identified by domain knowledge, are adjusted according to the predictions from the regression models. This step helps maintain feature consistency and reduces the likelihood of generating detectable anomalies.